OpenStudy (anonymous):

(152-10y)/6=(96-6y)/4 I need help showing the steps and answering this question

OpenStudy (anonymous):

Multiply both sides by 6 or 4..

OpenStudy (tyteen4a03):

Make use of cross-multiplication policy, where a/b = c/d would be transformed into ad = bc.

OpenStudy (anonymous):